OpenStudy (anonymous):

Are the given vectors normal? A=(5,-2) and B=(6,15)

OpenStudy (irishboy123):

\[\vec a • \vec b = |\vec a|| \vec b|\cos \theta = a_xb_x + a_yb_y = 0\] if they are normal [because cos (pi/2) = 0]

OpenStudy (anonymous):

thank you

OpenStudy (irishboy123):

ie if vectors a & b are normal [because cos (pi/2) = 0]